# Zhaoxiang Zhang9, Yuanxing Zhang10, Ge Zhang3,7, Ke Xu1, Jie Fu11*, Wenhao Huang2,3

### v3

@@ Appendix C Supervised Fine-Tuning Details
Image QA Vision-Flan (Xu et al., 2023), VizWiz (Gurari et al., 2018), LAION-GPT4V [^6] https://huggingface.co/datasets/laion/gpt4v-dataset, LLaVAR (Zhang et al., 2023c), OCR-VQA (Mishra et al., 2019), VQA (Goyal et al., 2016), TextVQA (Singh et al., 2019), OK-VQA (Marino et al., 2019), Mantis-Instruct (Jiang et al., 2024)
@@
Video Generation MSR-VTT (Xu et al., 2016), MSVD (Chen & Dolan, 2011b)
Video Understanding MSR-VTT (Xu et al., 2016), MSVD (Chen & Dolan, 2011b), MSVD-QA (Chen & Dolan, 2011a), MSRVTT-QA (Xu et al., 2017)
@@
Similar to pre-training (c.f., Appendix B), we enable Flash Attention (Dao et al., 2022; Dao, 2023) during supervised fine-tuning. Gradient clipping is set to 1.0. The maximum sequence length for training is 2800 tokens. We use a cosine learning rate scheduler with a peak learning rate of 3e-5 and a warmup ratio of 0.03. The optimizer used is AdamW (Loshchilov & Hutter, 2017).
@@ E.1 Image Generation Evaluation
We compute two additional automatic metrics for evaluating image generation, i.e., SSIM (Wang et al., 2004) and Aesthetic Predictor v2.5777https://github.com/discus0434/aesthetic-predictor-v2-5?tab=readme-ov-file for the evaluation of structural integrity and aesthetics, respectively. SSIM (Structural Similarity Index Measure) evaluates the perceptual similarity between the generated images and the ground-truth images, focusing on luminance, contrast, and structure, with scores ranging from -1 (dissimilar) to 1 (identical). Aesthetic Predictor V2.5 is a SigLIP (Zhai et al., 2023)-based predictor that evaluates the aesthetics of an image on a scale from 1 to 10 (10 is the best). In addition, we randomly select 100 image descriptions from MS-COCO test set, and used each model to generate images accordingly for human preference evaluation. We ask 3 annotators to rank 3 images generated by the 3 models: “given the image description, which image is preferred?” The average ranking of MIO’s, AnyGPT’s, and Emu’s generated images are 1.2 (MIO), 2.9 (AnyGPT), 1.9 (Emu). MIO aligns the best with the human preference. The percentage agreement between the three annotators (calculated as the number of cases with identical rankings by all annotators divided by 100) is 82.3%, indicating a high consistency in the human evaluation.
@@ E.2 Speech-to-Speech Evaluation
E.3 TTS Evaluation
@@ E.4 Loss Curves
E.5 More Demonstrations.

### v4

@@ Appendix C Supervised Fine-Tuning Details
Image QA Vision-Flan (Xu et al., 2023), VizWiz (Gurari et al., 2018), LAION-GPT4V , LLaVAR (Zhang et al., 2023c), OCR-VQA (Mishra et al., 2019), VQA (Goyal et al., 2016), TextVQA (Singh et al., 2019), OK-VQA (Marino et al., 2019), Mantis-Instruct (Jiang et al., 2024)
@@
Video Generation MSR-VTT (Xu et al., 2016), MSVD (Chen and Dolan, 2011b)
Video Understanding MSR-VTT (Xu et al., 2016), MSVD (Chen and Dolan, 2011b), MSVD-QA (Chen and Dolan, 2011a), MSRVTT-QA (Xu et al., 2017)
@@
Similar to pre-training (c.f., Appendix B), we enable Flash Attention (Dao et al., 2022; Dao, 2023) during supervised fine-tuning. Gradient clipping is set to 1.0. The maximum sequence length for training is 2800 tokens. We use a cosine learning rate scheduler with a peak learning rate of 3e-5 and a warmup ratio of 0.03. The optimizer used is AdamW (Loshchilov and Hutter, 2017).
@@ E.1 Image Generation Evaluation
In Table 13, we compute two additional automatic metrics for evaluating image generation, i.e., SSIM (Wang et al., 2004) and Aesthetic Predictor v2.5555https://github.com/discus0434/aesthetic-predictor-v2-5?tab=readme-ov-file for the evaluation of structural integrity and aesthetics, respectively. SSIM (Structural Similarity Index Measure) evaluates the perceptual similarity between the generated images and the ground-truth images, focusing on luminance, contrast, and structure, with scores ranging from -1 (dissimilar) to 1 (identical). Aesthetic Predictor V2.5 is a SigLIP (Zhai et al., 2023)-based predictor that evaluates the aesthetics of an image on a scale from 1 to 10 (10 is the best). In addition, we randomly select 100 image descriptions from MS-COCO test set, and used each model to generate images accordingly for human preference evaluation. We ask 3 annotators to rank 3 images generated by the 3 models: “given the image description, which image is preferred?” The average ranking of MIO’s, AnyGPT’s, and Emu’s generated images are 1.2 (MIO), 2.9 (AnyGPT), 1.9 (Emu). MIO aligns the best with the human preference. The percentage agreement between the three annotators (calculated as the number of cases with identical rankings by all annotators divided by 100) is 82.3%, indicating a high consistency in the human evaluation.
@@ E.2 Speech-to-Speech Evaluation
MIO Win 54%
Tie 25%
MIO Lose 21%
Table 15: Human evaluation for the TTS performance.
@@
Figure 2: Loss curves of pretraing stages.
E.3 TTS Evaluation
@@ E.4 Loss Curves
E.5 Demonstrations.
We illustrate the basic and advanced abilities of MIO in Figure 5 and 4. The basic abilities of MIO involve image understanding and generation, video understanding and generation, ASR, and TTS. The advanced abilities of MIO are based on its any-to-any and multimodal interleaved sequence generation features. These abilities involve visual storytelling (i.e., interleaved video-text generation), chain of visual thought, speech-in-speech-out, instructional image editing, visual guideline generation, etc. Figure 6 shows more demonstrations including multimodal chain of thought and in-context learning.
@@
E.6 More Ablation Studies.
Effect of Different Image Tokenizers.
The image tokenizer has a significant impact on image modality alignment. In Figure 3, we compare the image generation performance under a controlled setting after training for 3K steps in Stage I, using various image tokenizers. The image tokenizers for comparison include a VQGAN (Esser et al., 2020) with a vocabulary size of 1024 (VQGAN-1024), as well as the VQGAN-Gumbel with a vocabulary size of 8192 (VQGAN-8192) [^7] https://github.com/CompVis/taming-transformers. Our results indicate that the SEED-Tokenizer, which captures more semantic and higher-level image information, exhibits faster convergence. In contrast, both VQGAN tokenizers show slower convergence due to their lower-level image information.